# Artificial Intelligence AUC Area Under the Curve

### v2

@@ Abstract
Initialization plays a critical role in Deep Neural Network training, directly influencing convergence, stability, and generalization. Common approaches such as Glorot and He initializations rely on randomness, which can produce uneven weight distributions across layer connections. In this paper, we introduce the Sinusoidal initialization, a novel deterministic method that employs sinusoidal functions to construct structured weight matrices expressly to improve the spread and balance of weights throughout the network while simultaneously fostering a more uniform, well-conditioned distribution of neuron activation states from the very first forward pass. Because Sinusoidal initialization begins with weights and activations that are already evenly and efficiently utilized, it delivers consistently faster convergence, greater training stability, and higher final accuracy across a wide range of models, including convolutional neural networks, vision transformers, and large language models. On average, our experiments show an increase of $4.9\%$ in final validation accuracy and $20.9\%$ in convergence speed. By replacing randomness with structure, this initialization provides a stronger and more reliable foundation for Deep Learning systems.
@@
3Qsimov Quantum Computing S.L.
11footnotetext: These authors contributed equally.
@@ 1 Introduction
Weight initialization remains a critical yet often underappreciated component in the successful training of Deep Neural Networks (including Convolutional Neural Networks and Transformer-based architectures). Although numerous training strategies and architectural innovations receive substantial attention, initialization itself frequently remains in the background, implicitly accepted as a solved or trivial problem. Nonetheless, the initial setting of weights profoundly affects training dynamics, influencing convergence speed, stability, and even the final model’s ability to generalize (LeCun et al., 1998; Glorot and Bengio, 2010; He et al., 2015b).
Historically, weight initialization has evolved significantly. Early methods such as those proposed by LeCun et al. (1998) were primarily heuristic, offering basic variance scaling principles. The seminal work of Glorot and Bengio (2010) rigorously connected weight initialization to variance preservation across layers, introducing a randomized scaling approach that became foundational for Deep Learning (DL). Later, He et al. (2015b) refined this idea, tailoring variance preservation specifically to networks with Rectified Linear Units, and their approach quickly established itself as a default choice in modern DNN training pipelines.
@@
In this paper, we directly address this unexplored assumption by proposing a novel deterministic initialization scheme, termed Sinusoidal initialization. This technique preserves the critical variance-scaling property central to the effectiveness of Glorot and He initialization schemes, while completely removing randomness from the initialization procedure. It achieves this by employing Sinusoidal functions to deterministically initialize weight matrices, ensuring that each neuron in a given layer receives a distinct weight configuration. This use of Sinusoidal patterns enables rich diversity in weight values while maintaining global variance characteristics, striking a balance between structure and expressive power.
@@
By challenging the ingrained belief in randomness as a necessary condition for initialization, we open new avenues for rethinking foundational aspects of DNN training, advocating for deterministic and comprehensive practices in the future of DL research.
@@ 2 Background: Revisiting Initialization
Effective weight initialization is deeply rooted in the principles of variance preservation, which aim to maintain stable signal propagation across layers. This stability is crucial to avoid the notorious problem of vanishing or exploding gradients, which hinder DNN training (Hochreiter et al., 2001).
Formally, for a layer with weight matrix $W\in\mathbb{R}^{m\times n}$, Glorot and Bengio (2010) proposed Glorot initialization, where each weight is sampled from a distribution (either uniform or normal) with zero mean and variance $\textup{Var}(W)=2/(m+n)$, specifically chosen to balance the flow of information during both the forward and backward passes.
Subsequently, Glorot et al. (2011) observed that ReLUs produce naturally sparse activations, typically leaving roughly $50\%$ of hidden units active (positive) and the other 50% inactive (zero). They highlighted the significance of this balanced activation pattern for efficient gradient flow and effective learning. Recognizing this characteristic, He et al. (2015b) refined this initialization specifically for networks using ReLU activations, proposing the variance $\textup{Var}(W)=2/n$ which analytically accounted for ensuring robust data propagation.
@@
Various methods have attempted to reduce or structure this randomness. Orthogonal initialization (Saxe et al., 2014), for instance, ensures orthogonality of weight matrices, stabilizing signal propagation and improving gradient norms, but still involves random orthogonal matrix sampling. Layer-Sequential Unit-Variance (LSUV) (Mishkin and Matas, 2016) takes a step further by adjusting randomly initialized weights iteratively until achieving a precise activation variance. However, LSUV remains data-dependent and partially stochastic, limiting direct reproducibility and practical convenience.
A notable gap thus remains in the literature: Despite these advances, there is currently no widely adopted deterministic initialization method that fully preserves the critical statistical properties of Glorot and He initializations without involving any stochastic component. This unexplored direction motivates our work, wherein we propose a fully deterministic alternative, explicitly designed to overcome these limitations.
@@ 3.1 Sinusoidal Initialization
The amplitude $a$ is chosen according to theoretical foundations established by Glorot and Bengio (2010) and He et al. (2015b), which dictate the variance of weights in order to enable efficient signal propagation. Specifically, in our particular case, we adhere to the variance proposed by Glorot, setting the variance of $W$ to be $2/(m+n)$. If $v$ denotes the variance of $W$ when $a=1$, it then follows that $2/(m+n)=a^{2}v$, which yields a natural way to determine the amplitude $a$.
@@
Unlike traditional random initializations, this Sinusoidal scheme yields a non-Gaussian distribution of weights. As illustrated in Figure 3, the weights follow an arcsine distribution, exhibiting higher density near the extrema ±$\pm a$ and lower density near zero. This contrasts with the symmetric bell-shaped curves centered at zero typically produced by standard stochastic initialization methods. Additionally, Figure 3 provides a heatmap of matrix $W$, showcasing the underlying harmonic structure induced by the Sinusoidal construction. The structured and oscillatory pattern of each row is reminiscent of sinusoidal positional encodings employed in Transformer architectures (Vaswani et al., 2017), thereby offering spatial diversity without relying on randomness.
@@ 3.2 Theoretical Motivation
Figure 4: Neuron activation states, in white active neurons ($>0$), for different initializations.
Traditional initializations exhibit distinct vertical white or black bands, indicating neurons that remain consistently active or inactive—i.e., skewed neurons. As noted in Section 2Glorot et al. (2011) emphasized the importance of balanced ReLU activations, with roughly $50\%$ of neurons active to ensure effective learning. Skewed neurons disrupt this balance, limiting the network’s expressive power by underutilizing its nonlinearity and pushing it toward more linear behavior. In contrast, the Sinusoidal initialization yields a more irregular, noise-like pattern with no dominant columns, suggesting a more balanced and expressive activation regime.
@@
This quantity captures the imbalance between positive and negative weights. Despite the zero-mean assumption, the value of $S$ can deviate substantially from zero due to statistical fluctuation. For example, under the He initialization (He et al., 2015b), where weights are generated with variance $\sigma^{2}=2/n$, the Central Limit Theorem implies that $S\sim\mathcal{N}(0,2)$.
@@
This theoretical result tightly links the statistic $S$ with the skewness property of neurons. In the particular case of the Sinusoidal initialization, this connection becomes especially meaningful. Unlike standard random initialization schemes which offer no guarantees on the aggregate behavior of individual neurons, the Sinusoidal initialization exhibits an inherent structural regularity. This symmetry manifests in the exact cancellation of weights along each row of the initialization matrix, ensuring that each neuron begins with a perfectly balanced contribution across its inputs. The following result makes this property precise:
@@
Let $W\in\mathbb{R}^{m\times n}$ be defined as in Equation (1). Then, for every row index $i\in\{1,2,\dots,m\}$, the entries along that row sum to zero:
@@
Consequently, if the objective is to initialize the weights without introducing any biased (i.e., skewed) neurons, then, at least in the first layer, we should aim to enforce $S=0$ for every neuron.
@@
In order to further look into more intrinsic horizontal dependencies, the Overfitting–Underfitting Indicator (OUI), as introduced by the authors in  Fernández-Hernández et al. (2025), can be employed to numerically assess functional independence across neurons. This indicator measures the similarity between activation patterns of sample pairs within a given layer, assigning values in the range $[0,1]$, where 0 indicates highly correlated activations and 1 denotes complete independence. In this context, the activation pattern of a sample is understood as a binary vector, where each entry corresponds to the on/off activation state of a hidden neuron. Hence, OUI quantifies how differently pairs of samples activate the network, providing a direct measure of expressive capacity at initialization.
While all methods yield reasonably high OUI scores—Glorot $(0.56)$, He $(0.59)$, Orthogonal $(0.57)$, and LSUV $(0.58)$—the Sinusoidal initialization consistently achieves the highest value, reaching an outstanding $0.98$. This indicates that, although traditional initializations already promote a significant degree of activation diversity, the Sinusoidal scheme provides a strictly better starting point with respect to this metric, offering a more expressive and balanced activation regime from the very beginning.
@@ 4 Experimental Benchmarking of Sinusoidal initialization
Our experimental validation specifically addresses two main hypotheses: 1) Accuracy improvement: The Sinusoidal initialization achieves higher final validation accuracy compared to alternative initialization schemes in a significant proportion of cases. 2) Accelerated convergence: The Sinusoidal initialization converges faster than any other initialization considered, as measured by the Area Under the Curve (AUC) metric, which besides for classification evaluation, can be used to compare the training convergence speeds (Huang and Ling, 2005).
To evaluate these hypotheses rigorously, we conducted experiments involving five diverse configurations of DNNs and datasets: ResNet-50 (He et al., 2015a) trained on CIFAR-100 Krizhevsky et al. (2009), MobileNetV3 (Howard et al., 2019) trained on Tiny-ImageNet Le and Yang (2015), EfficientNetV2 (Tan and Le, 2021) trained on Tiny-ImageNet, ViT-B16 (Dosovitskiy et al., 2021) trained on ImageNet-1K Russakovsky et al. (2015)), and BERT-mini (Devlin et al., 2019) trained on WikiText Merity et al. (2016). For the first three configurations, we utilized SGD, Adam, and AdamW optimizers, while ViT-B16 and BERT-mini were trained exclusively with SGD and AdamW, respectively, due to the inability of other optimizers to achieve effective training. To ensure a fair comparison between training processes, learning rate schedulers were not used, which may affect the comparability of final accuracies with benchmarks focused on maximizing accuracy.
We trained each configuration fully from scratch using four initialization methods: Default (variants of the He initialization detailed in Table 2), Orthogonal (Orth.), LSUV, and our proposed Sinusoidal initialization (Sin.). The exact training procedures, including hyperparameter settings and implementation details, are provided in Appendix C.3. We measured the maximum validation accuracy achieved at epoch 1, epoch 10, and overall, as well as the AUC for every training scenario. The detailed experimental outcomes are summarized in Table 3.
@@
Table 3: Accuracy (%) after 1 epoch, 10 epochs, and at convergence, as well the AUC for each optimizer and initialization scheme (Default, Orthogonal, LSUV, Sinusoidal).
Model / Dataset Optim. 1 epoch accuracy (%) 10 epoch accuracy (%) Maximum accuracy (%) AUC
@@
ViT-16 ImageNet-1k SGD $2.8$ $3.8$ $3.5$ $2.6$ $15.2$ $15.2$ $14.1$ $13.9$ $28.6$ $28.2$ $29.6$ $31.5$ $23.3$ $25.2$ $24.1$ $25.4$
BERT-mini WikiText AdamW $9.4$ $6.3$ $11.0$ $6.3$ $13.6$ $10.9$ $14.8$ $17.0$ $40.4$ $42.2$ $15.9$ $41.1$ $58.2$ $71.6$ $32.2$ $72.0$
@@
Analyzing these tables, several clear trends emerge. Firstly, our Sinusoidal initialization consistently achieves higher final validation accuracy compared to other initializations, supporting our first hypothesis. Notably, in the case of MobileNetV3 on TinyImageNet, Orthogonal initialization slightly outperforms our method. Despite this isolated exception, the general trend strongly favors our Sinusoidal initialization. This particular case, however, highlights that this research field remains rich with open avenues for exploration. The behavior observed in this architecture offers a valuable opportunity to further advance our understanding of the DNN training process and refine Sinusoidal initialization strategy accordingly.

### v3

@@ Abstract
Initialization plays a critical role in Deep Neural Network training, directly influencing convergence, stability, and generalization. Common approaches such as Glorot and He initializations rely on randomness, which can produce uneven weight distributions across layer connections. In this paper, we introduce the Sinusoidal initialization, a novel deterministic method that employs sinusoidal functions to construct structured weight matrices expressly to improve the spread and balance of weights throughout the network while simultaneously fostering a more uniform, well‑conditioned distribution of neuron activation states from the very first forward pass. Because Sinusoidal initialization begins with weights and activations that are already evenly and efficiently utilized, it delivers consistently faster convergence, greater training stability, and higher final accuracy across a wide range of models, including convolutional neural networks, vision transformers, and large language models. On average, our experiments show an increase of $4.9\%$ in final validation accuracy and $20.9\%$ in convergence speed. By replacing randomness with structure, this initialization provides a stronger and more reliable foundation for Deep Learning systems.
@@
3Openchip & Software Technologies
footnotetext: These authors contributed equally to this work.footnotetext: Code available at: https://github.com/jmiravet/Sinusoidal-Initialization
@@ 1 Introduction
Weight initialization remains a critical yet often underappreciated component in the successful training of Deep Neural Networks (including Convolutional Neural Networks and Transformer-based architectures). Although numerous training strategies and architectural innovations receive substantial attention, initialization itself frequently remains in the background, implicitly accepted as a solved or trivial problem. Nonetheless, the initial setting of weights profoundly affects training dynamics, influencing convergence speed, stability, and even the final model’s ability to generalize (lecun1998efficient; glorot2010understanding; he2015delving).
Historically, weight initialization has evolved significantly. Early methods such as those proposed by lecun1998efficient were primarily heuristic, offering basic variance scaling principles. The seminal work of glorot2010understanding rigorously connected weight initialization to variance preservation across layers, introducing a randomized scaling approach that became foundational for Deep Learning (DL). Later, he2015delving refined this idea, tailoring variance preservation specifically to networks with Rectified Linear Units, and their approach quickly established itself as a default choice in modern DNN training pipelines.
@@
In this paper, we directly address this scarcely explored assumption by proposing a novel deterministic initialization scheme, termed Sinusoidal initialization. This technique preserves the critical variance-scaling property central to the effectiveness of Glorot and He initialization schemes, while completely removing randomness from the initialization procedure. It achieves this by employing Sinusoidal functions to deterministically initialize weight matrices, ensuring that each neuron in a given layer receives a distinct weight configuration. This use of Sinusoidal patterns enables rich diversity in weight values while maintaining global variance characteristics, striking a balance between structure and expressive power.
@@
By challenging the ingrained belief in randomness as a necessary condition for initialization, we encourage new avenues for rethinking foundational aspects of DNN training, advocating for deterministic and comprehensive practices in the future of DL research.
@@ 2 Background: Revisiting Initialization
Effective weight initialization is deeply rooted in the principles of variance preservation, which aim to maintain stable signal propagation across layers. This stability is crucial to avoid the notorious problem of vanishing or exploding gradients, which hinder DNN training (hochreiter2001gradient).
Formally, for a layer with weight matrix $W\in\mathbb{R}^{m\times n}$, glorot2010understanding proposed Glorot initialization, where each weight is sampled from a distribution (either uniform or normal) with zero mean and variance $\textup{Var}(W)=2/(m+n)$, specifically chosen to balance the flow of information during both the forward and backward passes.
Subsequently, glorot2011deep observed that ReLUs produce naturally sparse activations, typically leaving roughly $50\%$ of hidden units active (positive) and the other 50% inactive (zero). They highlighted the significance of this balanced activation pattern for efficient gradient flow and effective learning. Recognizing this characteristic, he2015delving refined this initialization specifically for networks using ReLU activations, proposing the variance $\textup{Var}(W)=2/n$ which analytically accounted for ensuring robust data propagation.
@@
Various methods have attempted to reduce or structure this randomness. Orthogonal initialization (saxe2013exact), for instance, ensures orthogonality of weight matrices, stabilizing signal propagation and improving gradient norms, but still involves random orthogonal matrix sampling. Layer-Sequential Unit-Variance (LSUV) (mishkin2015all) takes a step further by adjusting randomly initialized weights iteratively until achieving a precise activation variance. However, LSUV remains data-dependent and partially stochastic, limiting direct reproducibility and practical convenience.
Building on this line of work, some recent approaches have explored fully deterministic alternatives to random initialization. blumenfeld2020beyond state that DNNs can be trained from nearly all-zero, symmetric initializations, as long as some form of symmetry-breaking, such as dropout or hardware-level non-determinism, is present. zhao2022zero propose a scheme based on identity and Hadamard transforms, which achieves benchmark performance on ResNet architectures while preserving signal propagation. While both methods demonstrate that randomness is not strictly necessary for successful training, they primarily aim to match the performance of standard random initializations in terms of final accuracy. Crucially, they do not report improvements in convergence speed, robustness across architectures, or theoretical insights into the limitations of randomness. These works play an important role in establishing deterministic initialization as a viable alternative, but remain early steps in the broader effort to understand and fully exploit its potential.
A notable gap thus remains in the literature: Despite these advances, there is currently no widely adopted deterministic initialization method with demonstrated potential to surpass the performance of random initializations. This scarcely explored direction motivates our work, wherein we propose a fully deterministic alternative, explicitly designed to deliver improved accuracy and convergence speed across a wide range of architectures.
@@ 3.1 Sinusoidal Initialization
The amplitude $a$ is chosen according to theoretical foundations established by glorot2010understanding and he2015delving, which dictate the variance of weights in order to enable efficient signal propagation. Specifically, in our particular case, we adhere to the variance proposed by Glorot, setting the variance of $W$ to be $2/(m+n)$. If $v$ denotes the variance of $W$ when $a=1$, it then follows that $2/(m+n)=a^{2}v$, which yields a natural way to determine the amplitude $a$.
@@
Unlike traditional random initializations, this Sinusoidal scheme yields a non-Gaussian distribution of weights. As illustrated in Figure 3, the weights follow an arcsine distribution, exhibiting higher density near the extrema $\pm a$ and lower density near zero. This contrasts with the symmetric bell-shaped curves centered at zero typically produced by standard stochastic initialization methods. Additionally, Figure 3 provides a heatmap of matrix $W$, showcasing the underlying harmonic structure induced by the Sinusoidal construction. The structured and oscillatory pattern of each row is reminiscent of sinusoidal positional encodings***Positional Encodings were indeed the initial inspiration behind our initialization. The reader may note some differences, such as the use of phase shifts in the waves, the exclusive use of sine (as opposed to both sine and cosine), and the adoption of increasing rather than decreasing frequencies. Nonetheless, the underlying connection lies in the expressive power of harmonic functions to assign independent vectors, a fact long understood since Fourier. employed in Transformer architectures (NIPS2017_3f5ee243), thereby offering spatial diversity without relying on randomness.
@@ 3.2 Theoretical Motivation
We develop the following theoretical analysis for DNNs employing ReLU-like activations, which are functions such as gelu (gelu), silu (silu), or prelu (prelu), whose shapes resemble the classical rectified activation. In such networks, we consider a neuron to be active if its output is positive, and inactive otherwise. Unlike ReLU, these activations may still propagate small negative values in the inactive state, resulting in a smoother transition between activation regimes.
@@
Figure 4: Neuron activation states, in white active neurons (>0>0), for different initializations.
Traditional initializations exhibit distinct vertical white or black bands, indicating neurons that remain consistently active or inactive—i.e., skewed neurons. As noted in Section 2glorot2011deep emphasized the importance of balanced ReLU activations, with roughly $50\%$ of neurons active to ensure effective learning. Skewed neurons disrupt this balance, limiting the network’s expressive power by underutilizing its nonlinearity and pushing it toward more linear behavior. In contrast, the Sinusoidal initialization yields a more irregular, noise-like pattern with no dominant columns, suggesting a more balanced and expressive activation regime.
@@
This quantity captures the imbalance between positive and negative weights. Despite the zero-mean assumption, the value of $S$ can deviate substantially from zero due to statistical fluctuation. For example, under the He initialization (he2015delving), where weights are generated with variance $\sigma^{2}=2/n$, the Central Limit Theorem implies that $S\sim\mathcal{N}(0,2)$.
@@
This theoretical result establishes a tight connection between the statistic $S$ and the skewness property of neurons. As a direct consequence, stochastic initializations obtained by sampling from any zero-mean distribution inevitably lead to skewed neurons with probability tending to one. Formally:
Theorem 2 (Asymptotic skewness under random initialization).
Consider a neuron with zero bias and input $X_{1},\ldots,X_{n}$ and weights $W_{1},\ldots,W_{n}$. Assume $\{W_{i}\}$ are i.i.d. with $\mathbb{E}[W_{1}]=0$ and $\operatorname{Var}(W_{1})=\theta^{2}\in(0,\infty)$, and $\{X_{i}\}$ are i.i.d., independent of $\{W_{i}\}$, with $\mathbb{E}[X_{1}]=\mu>0$ and $\operatorname{Var}(X_{1})=\sigma^{2}<\infty$. Let $Z_{n}=\sum_{i=1}^{n}W_{i}X_{i}$. Then for every deterministic sequence $\alpha_{n}\downarrow 0$, the probability of having
$\bigl|P(Z_{n}>0\mid W_{1},\ldots,W_{n})-1/2\bigr|>\alpha_{n}$
tends to one as $n\to\infty$. Equivalently, with probability tending to one there exists a nonzero skewness level at which the neuron is skewed.
Therefore, in order to design initializations that do not systematically induce skewed neurons, it is necessary to move beyond the classical family of random schemes that fall under the assumptions of Theorem 2. This motivates the exploration of deterministic initializations specifically crafted to enforce balance from the outset. In the particular case of the Sinusoidal initialization, the connection with Theorem 1 becomes especially meaningful. Unlike standard random initialization schemes which offer no guarantees on the aggregate behavior of individual neurons, the Sinusoidal initialization exhibits an inherent structural regularity. This symmetry manifests in the exact cancellation of weights along each row of the initialization matrix, ensuring that each neuron begins with a perfectly balanced contribution across its inputs. The following result makes this property precise:
@@
Let $W\in\mathbb{R}^{m\times n}$ be defined as in Equation (1) with $m<n$. Then, for every row index $i\in\{1,2,\dots,m\}$, the entries along that row sum to zero:
@@
Consequently, if the objective is to initialize the weights without introducing any biased (skewed) neurons, then, at least in the first layer, we should aim to enforce $S=0$ for every neuron.
@@
In order to further look into more intrinsic horizontal dependencies, the Overfitting–Underfitting Indicator (OUI), as introduced by the authors in  OUI, can be employed to numerically assess functional independence across neurons. This indicator measures the similarity between activation patterns of sample pairs within a given layer, assigning values in the range $[0,1]$, where 0 indicates highly correlated activations and 1 denotes complete independence. In this context, the activation pattern of a sample is understood as a binary vector, where each entry corresponds to the on/off activation state of a hidden neuron. Hence, OUI quantifies how differently pairs of samples activate the network, providing a direct measure of expressive capacity at initialization.
While all methods yield reasonably high OUI scores—Glorot $(0.56)$, He $(0.59)$, Orthogonal $(0.57)$, and LSUV (0.58)(0.58)—the Sinusoidal initialization consistently achieves the highest value, reaching an outstanding $0.98$. This indicates that, although traditional initializations already promote a significant degree of activation diversity, the Sinusoidal scheme provides a strictly better starting point with respect to this metric, offering a more expressive and balanced activation regime from the very beginning.
@@ 4 Experimental Benchmarking of Sinusoidal initialization
Our experimental validation specifically addresses two main hypotheses: 1) Accuracy improvement: The Sinusoidal initialization achieves higher final validation accuracy compared to alternative initialization schemes in a significant proportion of cases. 2) Accelerated convergence: The Sinusoidal initialization converges faster than any other initialization considered, as measured by the Area Under the Curve (AUC) metric, which besides for classification evaluation, can be used to compare the training convergence speeds (huang2005using).
To evaluate these hypotheses rigorously, we conducted experiments involving five diverse configurations of DNNs and datasets: ResNet-50 (he2015deepresiduallearningimage) trained on CIFAR-100 krizhevsky2009learning, MobileNetV3 (howard2019searchingmobilenetv3) trained on Tiny-ImageNet le2015tiny, EfficientNetV2 (tan2021efficientnetv2smallermodelsfaster) trained on Tiny-ImageNet, ViT-B16 (dosovitskiy2021imageworth16x16words) trained on ImageNet-1K ILSVRC15), and BERT-mini (devlin2019bertpretrainingdeepbidirectional) trained on WikiText wikitext. For the first three configurations, we utilized SGD, Adam, and AdamW optimizers, while ViT-B16 and BERT-mini were trained exclusively with SGD and AdamW, respectively, due to the inability of other optimizers to achieve effective training. To ensure a fair comparison between training processes, learning rate schedulers were not used, which may affect the comparability of final accuracies with benchmarks focused on maximizing accuracy.
We trained each configuration fully from scratch using four initialization methods: Default (variants of the He initialization detailed in Table 2), Orthogonal (Orth.), LSUV, and our proposed Sinusoidal initialization (Sin.). The exact training procedures, including hyperparameter settings and implementation details, are provided in Appendix C.3. We measured the maximum validation accuracy achieved at epoch 1, epoch 10, and overall, as well as the AUC for every training scenario.
@@
In all our setups, to fairly compare initializations using AUC across different runs, we ensure that all models are trained for the same number of epochs. For each combination of model, dataset, and optimizer, training is performed with early stopping to guarantee convergence for every initialization. Once the slowest run has converged, the remaining models continue training for the same number of additional epochs until reaching the maximum epoch count. This procedure ensures that (1) all trainings reach their maximum validation accuracy, and (2) the number of epochs used to calculate the AUC for measuring convergence speed is consistent across initializations. The detailed experimental outcomes are summarized in Table 3 and illustrated in Figure 6.
Table 3: Validation accuracy (%) after 1 epoch, 10 epochs, and maximum accuracy, as well the AUC for each optimizer and initialization scheme (Default, Orthogonal, LSUV, Sinusoidal).
Model /
Dataset
Optim. 1 epoch accuracy (%) 10 epoch accuracy (%) Maximum accuracy (%) AUC
@@
ViT‑16
ImageNet‑1k
SGD $2.8$ $3.8$ $3.5$ $2.6$ $15.2$ $15.2$ $14.1$ $13.9$ $28.6$ $28.2$ $29.6$ $31.5$ $23.3$ $25.2$ $24.1$ $25.4$
BERT‑mini
WikiText
AdamW $9.4$ $6.3$ $11.0$ $6.3$ $13.6$ $10.9$ $14.8$ $17.0$ $40.4$ $42.2$ $15.9$ $41.1$ $58.2$ $71.6$ $32.2$ $72.0$
@@
Analyzing these tables, several clear trends emerge. Firstly, our Sinusoidal initialization consistently achieves higher final validation accuracy compared to other initializations, supporting our first hypothesis. Notably, in the case of MobileNetV3 on TinyImageNet, Orthogonal initialization slightly outperforms our method. Despite this isolated exception, the general trend strongly favors our Sinusoidal initialization. This particular case, however, highlights that this research field remains rich with open avenues for exploration. We conducted an additional experiment replacing the HardSwish activation in MobileNetV3 with a ReLU, and under this setting, Sinusoidal once again outperformed all other initialization strategies, consistent with the rest of our results. This suggests that the observed behavior stems from the interaction between the HardSwish activation and the initialization scheme. This particular case, however, highlights that this research field remains rich with open avenues for exploration. The behavior observed in this architecture offers a valuable opportunity to further advance our understanding of the DNN training process and refine Sinusoidal initialization strategy accordingly.